# Identified-and-Targeted: The First Early Evidence of the Privacy-Invasive Use of Browser Fingerprinting for Online Tracking

### v1

@@ 4.2 Collecting Bids and HTTP Data
Figure 3: High level overview of cookie restoration experiment. In Step 1, FPTrace will control the browser to visit websites. In Step 2, FPTrace will export all cookies including including 1st party and 3rd party cookies. In A and A*, FPTrace will use the same browser fingerprints, and same experiment conditions. In B, FPTrace will use a different browser fingerprint. Then FPTrace will do cookie comparisons between A, A* and B.
4.3 Detecting Cookie Restoration
Our next step is to examine cookie restoration under various settings. In the "True Setting" ($S_{true}$), which is represented as A in Figure 3, we initiate a new browser profile without any pre-existing data, including browsing history and cookies. By using the OpenWPM framework to load the browser profile, and sequentially visit a list of websites, the profile is updated after each visit. The resulting profile is termed $P_{true}$, and the exported cookie from profile $P_{true}$ is designated as $C_{true}$.
@@ 4.3 Detecting Cookie Restoration
4.4 Detecting Fingerprinting Related Cookie Restoration
@@
4.5 GDPR and CCPA Configuration
  1. 0.
    New profile, True fingerprint.
  2. 1.
    Trained profile, True fingerprint, Have Cookies. (Base Line)
  3. 2.
    Trained profile, Fake fingerprint, Have Cookies.
  4. 3.
    Trained profile, True fingerprint, Have Cookies. (Same setting as 1)
  5. 4.
    Trained profile, Fake fingerprint, No Cookies.
  6. 5.
    Trained profile, True fingerprint, No Cookies.
Figure 4: High level overview of GDPR/CCPA experiment. Step A represents the training persona, with the original fingerprint. Step B involves Opt Out or Opt In actions on websites with ads, utilizing CMPs, and employing Prebid.js, or on the central Opt Out website NAI. Step C involves the collection of HTTP data. Step D represents the data analysis.
5 Experiment Setting
@@
5.1 Initial Profiling and Data Collection
The first step involves the creation of an initial profile by visiting a range of websites related to "Computers". The full website list of $W_{p}$ here is listed in Appendix A.2. During this phase, cookies are enabled, and various fingerprinting techniques are deployed to collect data for analysis. The recorded data includes fingerprinters used and data sharing activities. The generated profile is then saved for subsequent comparisons.
5.2 Variation in Scenarios
@@
    Profile without cookies is used to visit websites, allowing the collection of bids and cookies. (no cookies, true fingerprint, true IP).
@@
Table 1: Various configurations involving browser profiles, cookies, browser fingerprinting, and IP addresses.
Setting Number New Profile? Have Cookies? True Fingerprint? True IP?
a Yes No Yes Yes
b No Yes Yes Yes
c No Yes No Yes
d No No Yes Yes
e No No No Yes
5.3 Experiment Locations
@@
FPTrace was primarily implemented on devices situated in the United States, where we gathered bid data, monitored HTTP events, and conducted cookie restoration detection. In addition to examining regions lacking explicit privacy regulations in the United States, we also evaluated the use of browser fingerprinting in areas where privacy regulations are in place.
Privacy regulations like the General Data Protection Regulation (GDPR) [15] and the California Consumer Privacy Act (CCPA) [16] are designed to safeguard user personal information. Under GDPR, customers have assurance against privacy breaches until they consent on websites. For CCPA protection, customers must actively refuse consent on websites to avoid privacy breaches. In our experimental setup, we initially evaluate data sharing related to fingerprinting and cookie restoration using local IP addresses in locations within the United States, where privacy regulation protection was absent. Subsequently, we implement FPTrace on AWS servers located in Frankfurt, Germany, and California, USA. Frankfurt-based visits fall under GDPR protection, while those from California are under CCPA. The aim of selecting these varied regions is to discern the impact of browser fingerprinting in environments with and without privacy regulation protection.
For visitors accessing websites from Europe or California, USA, the sites are expected to present a consent panel, allowing them to either agree or refuse consent. CMPs are third-party services that facilitate this consent process for users. There is a variety of CMP services available. As identified by Liu et al.[33], four prominent CMPs - Cookiebot, Quantcast, Onetrust, and Didomi - are among the most widely used and can be considered for use in this context. FPTrace can automatically handle the opt-out and opt-in. To ensure compliance with Didomi, we first verify the existence of the consent dialog using Didomi.notice.isVisible. We then employ Didomi.setUserDisagreeToAll to opt out and proceed to conceal the consent dialog by adjusting the display attributes of its markup to ‘none’. Similarly, for OneTrust, we confirm the presence of the consent dialog via window.OneTrust, executing window.OneTrust.RejectAll to opt out and subsequently conceal the dialog. In the case of CookieBot, we validate the presence of the consent dialog through window.Cookiebot, navigating the DOM to locate the opt-out button with the ID CybotCookiebotDialogBodyButtonDecline and activating it. Regarding Quantcast, we search the DOM for the consent dialog using the qc-cmp2-summary-buttons class name, clicking the button labeled ‘Reject’ or similar. If the rejection option is not initially visible, we expand the dialog by selecting the ’more options’ button and then proceed to click ‘Reject All’. Additionally, we assess the effectiveness of the central opt-out service Network Advertising Initiative (NAI) [10], which enables users to refuse consent through a single visit to its website. Consequently, other websites should comply by not utilizing the users’ consent if this service is used.
6 Results and Evaluation
@@
In this section, we first present the results of our experiments conducted in the United States, where privacy regulation protection was absent. These experiments were focused on data sharing and cookie restoration through browser fingerprinting, which fall outside the scope of current privacy regulations. Following this, we examine whether data sharing based on browser fingerprinting also takes place under the GDPR and CCPA privacy regulations.
6.1 Bid Values and HTTP Events
6.1.1 Have Cookies
@@
We collect and analyze bids data across different settings, specifically transitioning from “have cookies, have data, true fingerprints, true IP address" to “have cookies, have data, fake fingerprints, true IP address." The sole differentiating factor here is the presence of “true/fake fingerprints." To ensure data consistency, we conduct the experiment setting “have cookies, have data, true fingerprints, true IP address" twice to verify minimal bid data and HTTP data fluctuations. Table 2 illustrates that bid values between two instances of “True FP True IP" are quite similar, meeting our objective of running them twice. However, when comparing “True FP True IP" to “Fake FP True IP," in Table 2, there are substantial differences in median and maximum bid values. This suggests that changes in browser fingerprinting impact bid values, implying that alterations in browser fingerprinting influence targeting and tracking in advertising.
@@
Table 2: The bid value in different fingerprints and IPs settings. Avg represents the average of all bid value. Median represents the median bid value. Min represents the minimum bid value. Max represents the maximum bid value. We can observe that the first row values are different from the other two rows, which are similar to each other, thus showing that the change in FP creates a more marked impact on the bid values indicating that fingeprinting is being used for tracking.
@@
We have also generated a CDF plot (see Figure 5) using bid data from the two experiment settings. The distributions of CDF between different settings are similar when the CDF score is lower than 0.8, as all bids values are lower than 1. So we only display the distributions when the CDF scroe is larger than 0.8. In this CDF plot, the trends for two instances of “True FP True IP" are also similar, signifying that bid data remains stable under the same browser fingerprint. However, the introduction of fake fingerprints alters the trend of bid data. This indicates a significant difference in the distribution of bids between the two settings. The CDF plot provides evidence that changes in browser fingerprinting affect targeting and tracking in advertising.
Figure 5: CDF of Different Fingerprints and IPs settings in Keeping Cookies. The range of bids value is from 0 to 5. The range of CDF score is between 0.8 to 1. We can observe that the red curve is much different from the other two curves which are closer to each other, thus showing that fingerprinting is being used for tracking.
Furthermore, we calculated the number of HTTP chains, syncing events, and total HTTP records using HTTP data from the two settings. The results are outlined in Table 3. It is evident that the number of total HTTP chains in two instances of “True FP True IP" is similar and exceeds 36,000. However, changing the browser fingerprint reduces this number to 6,345, a substantial drop. This reduction indicates fewer HTTP chains can be established, implying a decrease in redirects and a reduced ability for websites to recognize the browser profile. A similar pattern emerges in syncing events: the number of syncing events in “Fake FP True IP" is 50% less than in “True FP True IP," signifying a 50% reduction in data sharing. The total number of HTTP records in “Fake FP True IP" is nearly 65% less than in “True FP True IP." The decrease in HTTP records also implies reduced recognition on the internet.
Table 3: The number of events and data through HTTP data. Chains represents the number of chains we built using collected HTTP data. Syncing represents the number of chains containing the data sharing. Records represents the total number of HTTP records collected in this settings. We can observe that the first row values are different from the other two rows, which are similar to each other, thus showing that the change in FP creates a more marked impact on the bid values indicating that fingeprinting is being used for tracking.
@@
Figure 6: CDF of Different Fingerprints and IPs settings in Removing Cookies. The range of bids value is from 0 to 5. The range of CDF score is between 0.8 to 1. We can observe that the red curve is much different from the other two curves which are closer to each other, thus showing that fingerprinting is being used for tracking.
Takeaway [Have Cookies]: 1. Bids trend and range is changed after changing the browser fingerprinting. 2. HTTP chains, syncing events and total records drop after changing the browser fingerprinting.
6.1.2 No Cookies
We gather and scrutinize bid data across various scenarios, particularly transitioning from a state characterized by “no cookies, have data, true fingerprints, true IP address" to a state with “no cookies, have data, true fingerprints, true IP address". To ensure the consistency of the data, we repeat the experiment under the setting “have cookies, have data, true fingerprints, true IP address" to confirm minimal variations in bid data and HTTP data. In the context of “true fingerprints", we introduce one fabricated fingerprint, while “fake fingerprints" includes another counterfeit fingerprint. Table 4 provides an overview of bid values across these three different configurations. When we keep the fingerprints consistent and eliminate cookies, the average and median bid values experience marginal increases, but the maximum bid value rises significantly. Following the alteration of fingerprints, both the average and maximum bid values increase, while the median value experiences a notable decrease. This suggests that changes in browser fingerprinting have an impact on bid values, underscoring their influence on targeting and tracking in advertising.
Table 4: The bid value in different fingerprints and IPs settings. Avg represents the average of all bid value. Median represents the median bid value. Min represents the minimum bid value. Max represents the maximum bid value. Setting “True FP True IP 2” do keep the cookies. Other two settings do not keep the cookies.
@@ 6.1.2 No Cookies
We have also generated a CDF plot (see Figure 6) using bid data from the two experimental configurations. The distributions of CDF between different settings are similar when the CDF score is lower than 0.8, as all bids values are lower than 1. So we only display the distributions when the CDF scroe is larger than 0.8. In this CDF plot, the patterns for two instances of "True FP True IP" are notably similar, indicating that bid data remains stable when the same browser fingerprint is used. However, the introduction of fake fingerprints disrupts the trend of bid data, signifying a substantial difference in bid distribution between the two settings. The CDF plot furnishes evidence that variations in browser fingerprinting influence targeting and tracking in advertising.
@@
Table 5: The number of events and data through HTTP data. Chains represents the number of chains we built using collected HTTP data. Syncing represents the number of chains containing the data sharing. Records represents the total number of HTTP records collected in this settings. CR represents the setting of Cookie Removed. All data are from settings including removing cookies.
@@
Takeaway [No Cookies]: 1. Bids trend and range is changed after removing cookies and changing the browser fingerprinting. 2. HTTP chains, and total records drop after removing cookies and changing the browser fingerprinting.
@@
Based on our analysis of bid values and HTTP events, we can draw the conclusion that browser fingerprinting indeed plays a significant role in targeting and tracking within the realm of advertising.

### v2

@@ 4.2. Collecting Bids and HTTP Data
Figure 3. High level overview of cookie restoration experiment. In Step 1, FPTrace will control the browser to visit websites. In Step 2, FPTrace will export all cookies including including 1st party and 3rd party cookies. In A and A*, FPTrace will use the same browser fingerprints, and same experiment conditions. In B, FPTrace will use a different browser fingerprint. Then FPTrace will do cookie comparisons between A, A* and B.
4.3. Detecting Cookie Restoration
Our next step is to examine cookie restoration under various settings. In the ”True Setting” ($S_{true}$), which is represented as A in Figure 3, we initiate a new browser profile without any pre-existing data, including browsing history and cookies. By using the OpenWPM framework to load the browser profile, and sequentially visit a list of websites, the profile is updated after each visit. The resulting profile is termed $P_{true}$, and the exported cookie from profile $P_{true}$ is designated as $C_{true}$.
Table 1. Various configurations involving browser profiles, cookies, browser fingerprinting, and IP addresses.
Setting Number New Profile? Have Cookies? True Fingerprint? True IP?
a Yes No Yes Yes
b No Yes Yes Yes
c No Yes No Yes
d No No Yes Yes
e No No No Yes
@@ 4.3. Detecting Cookie Restoration
4.4. Detecting Fingerprinting Related Cookie Restoration
@@
5. Experiment Setting
@@
5.1. Initial Profiling and Data Collection
The first step involves the creation of an initial profile by visiting a range of websites related to ”Computers”. The full website list of $W_{p}$ here is listed in Appendix A.1. During this phase, cookies are enabled, and various fingerprinting techniques are deployed to collect data for analysis. The recorded data includes fingerprinters used and data sharing activities. The generated profile is then saved for subsequent comparisons.
5.2. Variation in Scenarios
@@
    Profile without cookies is used to visit websites, enabling bid and cookie collection (no cookies, true fingerprint, true IP).
@@
5.3. Experiment Locations
@@
FPTrace was primarily implemented on devices situated in the United States, where we gathered bid data, monitored HTTP events, and conducted cookie restoration detection. In addition to examining regions lacking explicit privacy regulations in the United States, we also evaluated the use of browser fingerprinting in areas where privacy regulations are in place. The location selection of GDPR and CCPA experiments are explained in Appendix A.4.
6. Results and Evaluation
@@
In this section, we first present the results of our experiments conducted in the United States, where privacy regulation protection was absent. These experiments were focused on data sharing and cookie restoration through browser fingerprinting, which fall outside the scope of current privacy regulations. Following this, we examine whether data sharing based on browser fingerprinting also takes place under the GDPR and CCPA privacy regulations in Appendix A.5 and A.7. The raw data we collected is listed on Onedrive (One, 2023).
6.1. Bid Values and HTTP Events
6.1.1. Have Cookies
@@
We collect and analyze bids data across different settings, specifically transitioning from “have cookies, have data, true fingerprints, true IP address” to “have cookies, have data, fake fingerprints, true IP address.” The sole differentiating factor here is the presence of “true/fake fingerprints.” To ensure data consistency, we conduct the experiment setting “have cookies, have data, true fingerprints, true IP address” twice to verify minimal bid data and HTTP data fluctuations. Table 2 illustrates that bid values between two instances of “True FP True IP” are quite similar, meeting our objective of running them twice. However, when comparing “True FP True IP” to “Fake FP True IP,” in Table 2, there are substantial differences in median and maximum bid values. This suggests that changes in browser fingerprinting impact bid values, implying that alterations in browser fingerprinting influence targeting and tracking in advertising.
Figure 4. Figure a is the CDF of different fingerprints and IPs settings in keeping cookies. Figrue b is the CDF of different fingerprints and IPs settings in removing cookies. The range of bids value is from 0 to 5. The range of CDF score is between 0.8 to 1. In Figure a, We can observe that the red curve is much different from the other two curves which are closer to each other, thus showing that fingerprinting is being used for tracking. In Figure b We can observe that the red curve is much different from the other two curves which are closer to each other, thus showing that fingerprinting is being used for tracking.
@@
Table 2. The bid value in different fingerprints and IPs settings. Avg represents the average of all bid value. Median represents the median bid value. Min represents the minimum bid value. Max represents the maximum bid value. We can observe that the first row values are different from the other two rows, which are similar to each other, thus showing that the change in FP creates a more marked impact on the bid values indicating that fingeprinting is being used for tracking.
@@
We have also generated a CDF plot (see Figure 4a) using bid data from the two experiment settings. The distributions of CDF between different settings are similar when the CDF score is lower than 0.8, as all bids values are lower than 1. So we only display the distributions when the CDF scroe is larger than 0.8. In this CDF plot, the trends for two instances of “True FP True IP” are also similar, signifying that bid data remains stable under the same browser fingerprint. However, the introduction of fake fingerprints alters the trend of bid data. This indicates a significant difference in the distribution of bids between the two settings. The CDF plot provides evidence that changes in browser fingerprinting affect targeting and tracking in advertising.
Furthermore, we calculated the number of HTTP chains, syncing events, and total HTTP records using HTTP data from the two settings. The results are outlined in Table 3. It is evident that the number of total HTTP chains in two instances of “True FP True IP” is similar and exceeds 36,000. However, changing the browser fingerprint reduces this number to 6,345, a substantial drop. This reduction indicates fewer HTTP chains can be established, implying a decrease in redirects and a reduced ability for websites to recognize the browser profile. A similar pattern emerges in syncing events: the number of syncing events in “Fake FP True IP” is 50% less than in “True FP True IP,” signifying a 50% reduction in data sharing. The total number of HTTP records in “Fake FP True IP” is nearly 65% less than in “True FP True IP.” The decrease in HTTP records also implies reduced recognition on the internet.
Table 3. The number of events and data through HTTP data. Chains represents the number of chains we built using collected HTTP data. Syncing represents the number of chains containing the data sharing. Records represents the total number of HTTP records collected in this settings. We can observe that the first row values are different from the other two rows, which are similar to each other, thus showing that the change in FP creates a more marked impact on the bid values indicating that fingeprinting is being used for tracking.
@@
Takeaway [Have Cookies]: (1) Bids trend and range are changed after changing the browser fingerprinting. (2) HTTP chains, syncing events and total records drop after changing the browser fingerprinting.
6.1.2. No Cookies
We gather and scrutinize bid data across various scenarios, particularly transitioning from a state characterized by “no cookies, have data, true fingerprints, true IP address” to a state with “no cookies, have data, true fingerprints, true IP address”. To ensure the consistency of the data, we repeat the experiment under the setting “have cookies, have data, true fingerprints, true IP address” to confirm minimal variations in bid data and HTTP data. In the context of “true fingerprints”, we introduce one fabricated fingerprint, while “fake fingerprints” includes another counterfeit fingerprint. Table 4 provides an overview of bid values across these three different configurations. When we keep the fingerprints consistent and eliminate cookies, the average and median bid values experience marginal increases, but the maximum bid value rises significantly. Following the alteration of fingerprints, both the average and maximum bid values increase, while the median value experiences a notable decrease. This suggests that changes in browser fingerprinting have an impact on bid values, underscoring their influence on targeting and tracking in advertising.
Table 4. The bid value in different fingerprints and IPs settings. Avg represents the average of all bid value. Median represents the median bid value. Min represents the minimum bid value. Max represents the maximum bid value. Setting “True FP True IP 2” do keep the cookies. Other two settings do not keep the cookies.
@@ 6.1.2. No Cookies
We have also generated a CDF plot (see Figure 4b) using bid data from the two experimental configurations. The distributions of CDF between different settings are similar when the CDF score is lower than 0.8, as all bids values are lower than 1. So we only display the distributions when the CDF scroe is larger than 0.8. In this CDF plot, the patterns for two instances of ”True FP True IP” are notably similar, indicating that bid data remains stable when the same browser fingerprint is used. However, the introduction of fake fingerprints disrupts the trend of bid data, signifying a substantial difference in bid distribution between the two settings. The CDF plot furnishes evidence that variations in browser fingerprinting influence targeting and tracking in advertising.
@@
Table 5. The number of events and data through HTTP data. Chains represents the number of chains we built using collected HTTP data. Syncing represents the number of chains containing the data sharing. Records represents the total number of HTTP records collected in this settings. CR represents the setting of Cookie Removed. All data are from settings including removing cookies.
@@
Takeaway [No Cookies]: (1) Bids trend and range are changed after removing cookies and changing the browser fingerprinting. (2) HTTP chains, and total records drop after removing cookies and changing the browser fingerprinting.
@@
Based on our analysis of bid values and HTTP events, we can draw the conclusion that browser fingerprinting plays a significant role in targeting and tracking within the realm of advertising.
6.2. Cookie Restoration
After the cookie data comparison, we totally detected 90 cookie key and host pairs. We documented 378 instances of cookie restoration related to fingerprinting across 90 unique combinations of cookie keys and host pairs across all settings. Subsequently, we conducted a comprehensive manual inspection of all 90 cookie key and host pair combinations. In light of our thorough experiments, we cannot definitively assert that browser fingerprinting is employed to restore cookies. The full details are listed in Appendix A.6.
7. Discussion and Limitations
Our experiment was conducted using IP addresses from two locations in the United States, both of which are located in the United States and are not subject to privacy regulations such as GDPR (GDP, 2016) or CCPA (CCP, 2018). In regions protected by such regulations, trackers like cookies are prohibited from tracking users once they opt out. However, our experiment has revealed that advertisers may employ browser fingerprinting to track users without providing any notification. It remains uncertain whether advertisers can continue using browser fingerprinting to track users, as there is currently no established framework for auditing advertisers in this context. It’s important to note that our experiment cannot be utilized to assess advertisers’ behavior within the constraints of privacy regulations.
Another limitation of our study is that all experiments were conducted on the Linux platform. We cannot determine whether users of Windows devices, MacOS devices, or mobile devices can still be tracked by advertisers using browser fingerprinting techniques. While some of our fake fingerprint data were obtained from Windows devices, MacOS devices, or mobile devices, which we used to emulate our experimental device browsers, it would be valuable to incorporate real Windows devices, MacOS devices, or mobile devices in the “True Fingerprints” settings to gain a more comprehensive understanding.
Compared to the work of Liu et al.(Liu et al., 2024), there are notable differences in our study in both experimental design and research objectives. While our work focuses on exploring various fingerprinting settings and assessing whether different privacy regulations can constrain fingerprinting techniques, Liu et al.(Liu et al., 2024) did not involve any specific fingerprinting configurations. Instead, their research aimed to evaluate whether CMPs, websites, or advertisers comply with users’ consent choices.
8. Related Work
Previous research has examined the connection between online tracking and ad targeting. Wills et al.(Wills and Tatar, 2012) explored Google and Facebook’s ad systems, identifying various ad types. Google sometimes showed non-contextual ads related to sensitive topics, even without relevant user activity. On Facebook, external browsing had no clear link to ads, but using the “Like” feature on third-party content influenced ad recommendations. Zeng et al.(Zeng et al., 2022) studied the impact of user attributes, demographics, and contextual factors on ad targeting and bid values. Using data from 286 participants across 10 websites, they found targeting was primarily shaped by the website, retargeting methods, and user behavior, with demographics playing a minor role. Variations in bid values were mostly linked to the website and user actions, highlighting the significance of contextual targeting. Cassel et al.(Cassel et al., 2021) analyzed web tracking across devices, noting fewer tracking requests on mobile. Privacy-focused browsers reduced tracking but showed susceptibility to fingerprinting, without delving into its role in ads. Fouad et al.(Fouad et al., 2022) investigated cookie restoration using VPNs, which might introduce latency. Our approach, using real IPs and manual inspections, revealed that factors like time and website data could lead to restoration resembling fingerprinting.
9. Conclusion
In this study, we address a gap in research and regulatory practices by investigating browser fingerprinting’s role in ad tracking. We introduce “FPTrace”, a framework for detecting fingerprinting usage, especially when changes in fingerprints disrupt ad targeting. FPTrace is evaluated in web environments where user data is routinely collected for ads. Our method involves A/B experiments, data leakage analysis, fingerprint manipulation, and ad bidding analysis to study how altered fingerprints affect cookie restoration, hypothesizing that prior user knowledge influences bid changes.
Key contributions include the FPTrace framework, integrated with OpenWPM, which simulates user interactions, collects bid data, and records HTTP data. Our findings show browser fingerprinting significantly impacts ad tracking, with noticeable differences in bid values and a reduction in HTTP records when fingerprints change. While fingerprints and cookies show some connection, evidence of direct involvement in cookie restoration remains inconclusive.
Acknowledgements.
We would like to thank anonymous reviewers for their helpful comments and feedback. This work was supported in part by National Science Foundation (NSF) under grants CCF-2317185, OAC-2139358, CNS- 2201465, and CNS-2154507.
Appendix A Appendix
A.1. Training Websites
  1. (1)
  2. (2)
  3. (3)
  4. (4)
  5. (5)
  6. (6)
  7. (7)
  8. (8)
  9. (9)
  10. (10)
  11. (11)
  12. (12)
  13. (13)
  14. (14)
  15. (15)
  16. (16)
  17. (17)
  18. (18)
  19. (19)
  20. (20)
  21. (21)
  22. (22)
  23. (23)
  24. (24)
  25. (25)
  26. (26)
  27. (27)
  28. (28)
  29. (29)
  30. (30)
  31. (31)
  32. (32)
  33. (33)
  34. (34)
  35. (35)
  36. (36)
  37. (37)
  38. (38)
  39. (39)
  40. (40)
A.2. Testing Websites
A.3. GDPR and CCPA Configuration
  1. (0)
    New profile, True fingerprint.
  2. (1)
    Trained profile, True fingerprint, Have Cookies. (Base Line)
  3. (2)
    Trained profile, Fake fingerprint, Have Cookies.
  4. (3)
    Trained profile, True fingerprint, Have Cookies. (Same setting as 1)
  5. (4)
    Trained profile, Fake fingerprint, No Cookies.
  6. (5)
    Trained profile, True fingerprint, No Cookies.
Figure 5. High level overview of GDPR/CCPA experiment. Step A represents the training persona, with the original fingerprint. Step B involves Opt Out or Opt In actions on websites with ads, utilizing CMPs, and employing Prebid.js, or on the central Opt Out website NAI. Step C involves the collection of HTTP data. Step D represents the data analysis.
A.4. Experiment Locations
Privacy regulations like the General Data Protection Regulation (GDPR) (GDP, 2016) and the California Consumer Privacy Act (CCPA) (CCP, 2018) are designed to safeguard user personal information. Under GDPR, customers have assurance against privacy breaches until they consent on websites. For CCPA protection, customers must actively refuse consent on websites to avoid privacy breaches. In our experimental setup, we initially evaluate data sharing related to fingerprinting and cookie restoration using local IP addresses in locations within the United States, where privacy regulation protection was absent. Subsequently, we implement FPTrace on AWS servers located in Frankfurt, Germany, and California, USA. Frankfurt-based visits fall under GDPR protection, while those from California are under CCPA. The aim of selecting these varied regions is to discern the impact of browser fingerprinting in environments with and without privacy regulation protection.
For visitors accessing websites from Europe or California, USA, the sites are expected to present a consent panel, allowing them to either agree or refuse consent. CMPs are third-party services that facilitate this consent process for users. There is a variety of CMP services available. As identified by Liu et al.(Liu et al., 2024), four prominent CMPs - Cookiebot, Quantcast, Onetrust, and Didomi - are among the most widely used and can be considered for use in this context. FPTrace can automatically handle the opt-out and opt-in. To ensure compliance with Didomi, we first verify the existence of the consent dialog using Didomi.notice.isVisible. We then employ Didomi.setUserDisagreeToAll to opt out and proceed to conceal the consent dialog by adjusting the display attributes of its markup to ‘none’. Similarly, for OneTrust, we confirm the presence of the consent dialog via window.OneTrust, executing window.OneTrust.RejectAll to opt out and subsequently conceal the dialog. In the case of CookieBot, we validate the presence of the consent dialog through window.Cookiebot, navigating the DOM to locate the opt-out button with the ID CybotCookiebotDialogBodyButtonDecline and activating it. Regarding Quantcast, we search the DOM for the consent dialog using the qc-cmp2-summary-buttons class name, clicking the button labeled ‘Reject’ or similar. If the rejection option is not initially visible, we expand the dialog by selecting the ’more options’ button and then proceed to click ‘Reject All’. Additionally, we assess the effectiveness of the central opt-out service Network Advertising Initiative (NAI) (nai, 2023), which enables users to refuse consent through a single visit to its website. Consequently, other websites should comply by not utilizing the users’ consent if this service is used.